How many litres of fuel costing $ 1.15 per litre must be mixed with 20 litres of fuel costing $ 0.85 per litre to obtain a mixture costing $ 1 per litre.
Answer by Maths68(1474)

Fuel one Amount = x
Fuel one Cost = $1.15 per liters
Fuel two Amount = 20
Fuel two Cost = $0.85 per liters
Mixture Amount= 20+x
Mixture cost = $1 per liters

(Fuel one Amount)* (Fuel one Cost) + (Fuel two Amount) * (Fuel two Cost) = Mixture Amount * Mixture cost
x*1.15 + 20*0.85 = (20+x)*1
1.15x+17=20+x
1.15x-x=20-17
0.15x=3
x=3/0.15
x=20

Fuel one Amount = x = 20 liters
Fuel one Cost = $1.15 per liters
Fuel two Amount = 20
Fuel two Cost = $0.85 per liters
Mixture Amount= 20+20 = 20+6 = 40 liters
Misture cost = $1 per liters


20 litres of fuel costing $ 1.15 per litre must be mixed with 20 litres of fuel costing $ 0.85 per litre to obtain a mixture of 40 liters fuel costing $ 1 per litre.